Find the area of a triangle bounded by the lines y=x, y=-x, and y=6?
1
Expert's answer
2012-10-30T09:59:42-0400
we have vertices
(0, 0) intersection y=x, y=-x
(6,6) intersection y=x, y=6
(-6,6) intersection y=6, y=-x
the height of this triangle is 6 inches and side oftriangle for this height is 12 inches so area=0.5*6*12=36 inches^2